OpenStudy (anonymous):

(3x10^2)(3x10^3)... scientific notation

OpenStudy (anonymous):

9E5

OpenStudy (anonymous):

??

OpenStudy (anonymous):

9E5 is the same as 9x10^5

OpenStudy (anonymous):

9x10^5

OpenStudy (anonymous):

not the correct answer...

OpenStudy (anonymous):

why???/////

OpenStudy (anonymous):

wrong symbol... my bad

OpenStudy (anonymous):

heyy just explain lolzzzzzzz

OpenStudy (anonymous):

\[3\times 10^2\times 3\times 10^3=3\times 3\times 10^2\times 10^3=9\times 10^5\]

OpenStudy (anonymous):

already got it thanks

OpenStudy (anonymous):

first equal sign by commutative law, last one by laws of exponents